OpenStudy (anonymous):

Find the solution to the equation 64^(2 – x) = 4^4x

jhonyy9 (jhonyy9):

64=4 on exponent 3 so than 3(2-x=4x 6-3x=4x 6=7x x=6/7

jhonyy9 (jhonyy9):

so thara is 3(2-x)=4x

OpenStudy (anonymous):

thank you

jhonyy9 (jhonyy9):

yw

OpenStudy (mertsj):